OpenStudy (amtran_bus):

Convert to logarithmic form -2^x+3

OpenStudy (amtran_bus):



jhonyy9 (jhonyy9):

so 3 is exponent too ?

OpenStudy (anonymous):

Start by subtracting the 3. -2^x = -3

OpenStudy (anonymous):

@qpHalcy0n there is no equation here, though.

OpenStudy (anonymous):

Then what can you really do here? There is no problem. Problem solved.

OpenStudy (amtran_bus):

So you cant convert it to a log? I didn't think you could have a log base -2 anyway.

OpenStudy (anonymous):

You didn't give enough information for us to do anything.

OpenStudy (amtran_bus):

See, I thought that since this was in exponential form, you could make it a log. I know it says +3, but that means it is moved up three on the graph.

OpenStudy (amtran_bus):

So how would you find the inverse of this problem?

OpenStudy (anonymous):

\[\text{So then you meant }y=-2^x+3\text{?}\\ \ \ \ y=-2^x+3\\ \ \ \ y-3=-2^x\\ \ \ \ 2^x=3-y\\ \ \ \ x=log_2[3-y] \]

OpenStudy (amtran_bus):

Thanks! Sorry, that is what I meant.